# Pareto Set Learning for Multi-Objective Reinforcement Learning

### v1

@@ A.1 Instantiation: PSL-MORL with DDQN
In this subsection, we introduce an instantiation of PSL-MORL by taking DDQN [Van Hasselt et al., 2016] as the single-objective RL algorithm. The pseudo code is shown in Algorithm 2.
PSL-MORL with DDQN first initializes the hypernetwork and HER. The parallel child processes are the same as that in Algorithm 1. After that, it operates on a batch of preferences. For each preference, the multi-dimensional interpolator $I(\bm{\omega})$ is used to obtain the interpolated preference $\bm{w}_{p}$, and a modified Q-learning update scheme (with the cosine similarity term $S_{c}$ proposed in [Basaklar et al., 2023]) is used to calculate the target value $\bm{y}$. The empirical expected loss is then obtained by calculating the DDQN loss for each preference. Finally, an optimizer (e.g., SGD) is employed to update the parameters of both the hypernetwork and target network.
@@ A.4 Benchmark Details
  • Fruit Tree Navigation (FTN): An innovative MORL benchmark introduced by  [Yang et al., 2019]. It involves a binary tree of depth $d$, where each leaf node is assigned a reward vector $\bm{r}\in\mathbb{R}^{6}$ representing the nutritional values of fruits, categorized into six types: Protein, Carbs, Fats, Vitamins, Minerals and Water. The goal of the agent is to navigate the tree to harvest fruit that optimizes these nutritional values based on specific preferences.
@@ A.5 Baselines Details
  • PD-MORL [Basaklar et al., 2023]: PD-MORL utilizes the preference as guidance to update the network parameters and covers the entire preference space.
@@ B.1 Preliminary
(Rademacher Complexity, Definition 13.1 in [Golowich et al., 2018]) Given a real-valued function class $\mathcal{H}$ and some set of data points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathcal{X}$, we define the (empirical) Rademacher complexity ^$\hat{\mathcal{R}}_{n}(\mathcal{H})$ as
@@ B.2 Rademacher Complexity of Neural Networks
In this section, we present an upper bound and a lower bound on the Rademacher complexity, given by  [Golowich et al., 2018], for the class of neural networks with parameter matrices of bounded Schatten norms. They will be used in our proof.
@@
[Golowich et al., 2018] Let $\mathcal{H}_{d}$ be the class of real-valued networks of depth $d$ over the domain $\mathcal{X}$, where each parameter matrix $W_{j}$ has Frobenius norm at most $M_{F}(j)$, and the activation function $\sigma$ is a $1$-Lipschitz, positive-homogeneous activation function which is applied element-wise (such as the ReLU). Then,
@@
[Golowich et al., 2018] Let $\mathcal{H}_{d}$ be the class of depth-$d$ neural networks over the domain $\mathcal{X}$, where each parameter matrix $W_{j}$ satisfies $\|W_{j}\|_{p}\leq M_{p}(j)$ for some Schatten $p$-norm $\|\cdot\|_{p}$, i.e., the parameter matrices $W_{1},\ldots,W_{d}$ in the $d$ layers have Schatten norms $\|\cdot\|_{p}$ upper-bounded by $M_{p}(1),\ldots,M_{p}(d)$, respectively. Then there exists a $\frac{1}{\gamma}$-Lipschitz loss $\ell$ and data points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathcal{X}$, such that

### v2

@@ A.1 Instantiation: PSL-MORL with DDQN
In this subsection, we introduce an instantiation of PSL-MORL by taking DDQN [Van Hasselt, Guez, and Silver, 2016] as the single-objective RL algorithm. The pseudo code is shown in Algorithm 2.
PSL-MORL with DDQN first initializes the hypernetwork and HER. The parallel child processes are the same as that in Algorithm 1. After that, it operates on a batch of preferences. For each preference, the multi-dimensional interpolator $I(\bm{\omega})$ is used to obtain the interpolated preference $\bm{w}_{p}$, and a modified Q-learning update scheme (with the cosine similarity term $S_{c}$ proposed in [Basaklar, Gumussoy, and Ogras, 2023]) is used to calculate the target value $\bm{y}$. The empirical expected loss is then obtained by calculating the DDQN loss for each preference. Finally, an optimizer (e.g., SGD) is employed to update the parameters of both the hypernetwork and target network.
@@ A.4 Benchmark Details
  • Fruit Tree Navigation (FTN): An innovative MORL benchmark introduced by  [Yang, Sun, and Narasimhan, 2019]. It involves a binary tree of depth $d$, where each leaf node is assigned a reward vector $\bm{r}\in\mathbb{R}^{6}$ representing the nutritional values of fruits, categorized into six types: Protein, Carbs, Fats, Vitamins, Minerals and Water. The goal of the agent is to navigate the tree to harvest fruit that optimizes these nutritional values based on specific preferences.
@@ A.5 Baselines Details
  • PD-MORL [Basaklar, Gumussoy, and Ogras, 2023]: PD-MORL utilizes the preference as guidance to update the network parameters and covers the entire preference space.
@@ B.1 Preliminary
(Rademacher Complexity, Definition 13.1 in [Golowich, Rakhlin, and Shamir, 2018]) Given a real-valued function class $\mathcal{H}$ and some set of data points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathcal{X}$, we define the (empirical) Rademacher complexity ^$\hat{\mathcal{R}}_{n}(\mathcal{H})$ as
@@ B.2 Rademacher Complexity of Neural Networks
In this section, we present an upper bound and a lower bound on the Rademacher complexity, given by  [Golowich, Rakhlin, and Shamir, 2018], for the class of neural networks with parameter matrices of bounded Schatten norms. They will be used in our proof.
@@
[Golowich, Rakhlin, and Shamir, 2018] Let $\mathcal{H}_{d}$ be the class of real-valued networks of depth $d$ over the domain $\mathcal{X}$, where each parameter matrix $W_{j}$ has Frobenius norm at most $M_{F}(j)$, and the activation function $\sigma$ is a $1$-Lipschitz, positive-homogeneous activation function which is applied element-wise (such as the ReLU). Then,
@@
[Golowich, Rakhlin, and Shamir, 2018] Let $\mathcal{H}_{d}$ be the class of depth-$d$ neural networks over the domain $\mathcal{X}$, where each parameter matrix $W_{j}$ satisfies $\|W_{j}\|_{p}\leq M_{p}(j)$ for some Schatten $p$-norm $\|\cdot\|_{p}$, i.e., the parameter matrices $W_{1},\ldots,W_{d}$ in the $d$ layers have Schatten norms $\|\cdot\|_{p}$ upper-bounded by $M_{p}(1),\ldots,M_{p}(d)$, respectively. Then there exists a $\frac{1}{\gamma}$-Lipschitz loss $\ell$ and data points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathcal{X}$, such that